OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 90 miles. After five hours the car has traveled 225 miles. Assuming a linear equation fits this situation, what are the ordered pairs that you would use to find the linear equation?

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

not funny miranda

OpenStudy (anonymous):

2h=90 5h=225

OpenStudy (anonymous):

thats not what im lookin forr

OpenStudy (anonymous):

Are you trying to find the cars velocity? :P

OpenStudy (anonymous):

no,ttrying to find a orderd pair thts a part of the linear equation

OpenStudy (anonymous):

tht fixes problem

OpenStudy (anonymous):

let m be the slope and b be the starting y then we would get 2 equations: 90=2*m+b 225=5*m+b multiplying the first equation with -1 and adding it to the second one we get: 135=3*m 135/3=m sub that in the first equation we get: 90=2*135/3+b 0=b so the linear equation would be y=135/3*m testing with both ordered pairs (2,90) (5,225): 90=135/3*2 correct 225=135/3*5 correct

OpenStudy (anonymous):

oh, yes, just searching for the ordered pairs... use (2,90) and (5,225) or, assuming the journey starts at 0 hours and 0 miles, use (0,0) and (2,90)